# B. Kaan Karamete Eli Glaser Kinetica DB Inc. 901 North Glebe Road, Arlington, Virginia 22203

### v1

@@ 5 Stochastic Gradient Descent
The next step is taking the derivative of the nodal average of the loss per each of these weight parameters and move against the direction of the gradient of each weight to minimize the loss. The incremental update on each unknown weigth is immediately made to reflect its impact on the next unknown weight variable computation as shown in Equation 4 and 5 in which $w_{j}^{(k+1)}$ is the next $(k+1)^{th}$ iteration on the $j^{th}$ weight parameter. This approach is not a guaranteed outcome in accelerating the convergence. Other alternative approaches such as batching or mini-batching discussed extensively in [13] might provide better computational outcome. However, we think this is beyond the scope of this study and our findings are satisfactory computationally for the cases we have tried so far. The iterations are continued if the relative incremental iteration delta of all unknown weights go below a preset user threshold (default value is 0.001) or the number of epoch iterations reaches the upper limit (default is 100) which is also a user prescribed parameter of the solver as shown in Figure 12. The convergence history plot between the number of iterations and the error is also shown in Figure 13 for the knowledge graphs of varying sizes from a few hundred to 100 million nodes. The trend in all cases is with early unstable fluctuations and rapid descending to the optimal as expected. The initial weight values and the rate of iteration, also known as training or learning rate, namely, $\beta$ as shown in Equation 4, can both also be overridden by the user.
@@
$\frac{\partial\sum\limits_{i}(Loss_{i})}{\partial w_{j}}=4\sum\limits_{i}\sum\limits_{m}(<f_{i},f_{m}>~{}\times~{}w_{j}s^{i}_{j}s^{m}_{j})$ (5)
@@ 6 Discussion and Conclusions
We have tried to map unlimited dimension general knowledge graph topology onto a $1-$dimensional vector embeddings by constructing the vector space from features that we think would best resemble local affinity and remote structure so that any vector similarity (inner product) between a node pair would result in the same similarity behavior if we had computed the Jaccard score with the number of common labels between the two nodes of every pair. To this end, the sub-features are chosen to be the predicates of hop-pattern numbers, cluster indices (computed by the recursive spectral bisection (RSB)), associated label indices, and the transitional probability (or the Pagerank score if weights are uniform) as explained in Section 2 above. The impact of the vector component sub-features on the similarity can be found by adding a weight parameter to multiply each of the sub-feature elements within the respective vector sub-range and test the result against an estimated ground truth as explained in Section 4. We have formulated the difference between the inner product of the assumed embeddings and the combined common-labels and Jaccard score as the ground truth, as our ad-hoc Loss function as shown in Equation 3. We then tried to optimize the nodal average of the total loss by applying a stochastic gradient descent (SGD) algorithm to find the unknown weights so that this total average loss is minimized as explained in Section 5 and Equations 4 and 5.

### v2

@@ 5 Stochastic Gradient Descent
The next step is taking the derivative of the nodal average of the loss per each of these weight parameters and move against the direction of the gradient of each weight to minimize the loss. The incremental update on each unknown weigth is immediately made to reflect its impact on the next unknown weight variable computation as shown in Equation 4 and 6 in which $w_{j}^{(k+1)}$ is the next $(k+1)^{th}$ iteration on the $j^{th}$ weight parameter. This approach is not a guaranteed outcome in accelerating the convergence. Other alternative approaches such as batching or mini-batching discussed extensively in [13] might provide better computational outcome. However, we think this is beyond the scope of this study and our findings are satisfactory computationally for the cases we have tried so far. The iterations are continued if the relative incremental iteration delta of all unknown weights go below a preset user threshold (default value is 0.001) or the number of epoch iterations reaches the upper limit (default is 100) which is also a user prescribed parameter of the solver as shown in Figure 12. The convergence history plot between the number of iterations and the error is also shown in Figure 13 for the knowledge graphs of varying sizes from a few hundred to 100 million nodes. The trend in all cases is with early unstable fluctuations and rapid descending to the optimal as expected. The initial weight values and the rate of iteration, also known as training or learning rate, namely, $\beta$ as shown in Equation 4, can both also be overridden by the user.
@@
$K_{i,m}=<f_{i},f_{m}>-\biggl{(}\alpha~{}jac(i,m)+$ (5)
$(1-\alpha)\frac{labels(i)\cap labels(m)}{\sum_{m}{distinct(labels(m))}}\biggl{)}$
$\frac{\partial\sum\limits_{i}(Loss_{i})}{\partial w_{j}}=4\sum\limits_{i}\sum\limits_{m}(K_{i,m}<f_{i},f_{m}>w_{j}s^{i}_{j}s^{m}_{j})$ (6)
@@ 6 Discussion and Conclusions
We have tried to map unlimited dimension general knowledge graph topology onto a $1-$dimensional vector embeddings by constructing the vector space from features that we think would best resemble local affinity and remote structure so that any vector similarity (inner product) between a node pair would result in the same similarity behavior if we had computed the Jaccard score with the number of common labels between the two nodes of every pair. To this end, the sub-features are chosen to be the predicates of hop-pattern numbers, cluster indices (computed by the recursive spectral bisection (RSB)), associated label indices, and the transitional probability (or the Pagerank score if weights are uniform) as explained in Section 2 above. The impact of the vector component sub-features on the similarity can be found by adding a weight parameter to multiply each of the sub-feature elements within the respective vector sub-range and test the result against an estimated ground truth as explained in Section 4. We have formulated the difference between the inner product of the assumed embeddings and the combined common-labels and Jaccard score as the ground truth, as our ad-hoc Loss function as shown in Equation 3. We then tried to optimize the nodal average of the total loss by applying a stochastic gradient descent (SGD) algorithm to find the unknown weights so that this total average loss is minimized as explained in Section 5 and Equations 4 and 6.